Fresh / Soph / Jr / Sr / Grad (circle one) College and Major:

The grade I expect to receive is:

Mid-Semester Evaluation

36-402: Advanced Data Analysis II February 25, 1998

• This evaluation form has only one purpose: to improve the quality of instruction during the remaining weeks of the semester.

• Below are several statements about the course, classroom, books, computer labs, etc. Please circle your response to each statement using the 5-point scale indicated.

• You should use the blank space after each statement (or on the back of each page) to offer further comments if just circling a number doesn’t adequately express your feelings.

• Your frank comments and constructive criticism would be greatly appreciated.

Do not sign your name.

Overall

My understanding of the material covered to date:

1  2  3  4  5

I’m totally lost Challenging, but I think I’m understanding the material This course is really easy

My interest in, and enjoyment of, this course to date:

1  2  3  4  5

I’m bored stiff It’s OK but I like another course better One of my favorite classes

Course

1.

The course is not too hard and not too easy.

Disagree  Agree
Strongly  Strongly
1 2 3 4 5

If you disagree, is it too \underline{hard} or too \underline{easy}? . . . . . . . . . . . . . . . . . . . .

2. The course is not too slow and not too fast.  1 2 3 4 5

If you disagree, is it too \underline{slow} or too \underline{fast}? . . . . . . . . . . . . . . . . . . . . .

3.

The things I am learning in this course will be useful in other classes or outside of school.

36-402: Advanced Data Analysis II Mid-Semester Evaluation

1 2 3 4 5 Instructor

1.

The instructor presents information in a clear and wellorganized manner.

Disagree  Agree
Strongly  Strongly
1 2 3 4 5

2.

The instructor speaks clearly, loudly enough, and is easy to understand.

1 2 3 4 5

3. The instructor answers questions willingly.  1 2 3 4 5

4.

The instructor uses questions or asks for comments to develop student/teacher interaction.

1 2 3 4 5

5. The instructor encourages critical or original thinking.  1 2 3 4 5

Textbook and Handouts

1.

The Venables and Ripley book Modern Applied Statistics using Splus is pretty good; I get a lot out of reading it.

Disagree  Agree
Strongly  Strongly
1 2 3 4 5

2.

The Kraus and Olson book The Basics of S and Splus is pretty good; I get a lot out of reading it.

1 2 3 4 5

3.

The handouts in class help me to understand important concepts in the course.

1 2 3 4 5

Other Comments on Instructor, Textbook, Handouts Written Work

1.

I get constructive feedback on my written work.

Disagree  Agree
Strongly  Strongly
1 2 3 4 5

2. Homework is returned promptly.  1 2 3 4 5

Classroom

1.

The overhead projector / transparencies are easy to see and read in class.

Disagree  Agree
Strongly  Strongly
1 2 3 4 5

2. The writing on the chalkboard is legible and easy to read.  1 2 3 4 5

3.

We don’t always do the same things in class; it doesn’t get boring.

1 2 3 4 5

Computer Demos and Presentations

1.

The in-class demo’s in Splus are helpful; I’d like to continue seeing them in the remaining weeks of the semester.

Disagree  Agree
Strongly  Strongly
1 2 3 4 5

2.

The text and graphics projected from the laptop are legible and easy to read.

1 2 3 4 5

Other Comments on Homework, Classroom, Computer Demos